The rate constant for the equation
2 C2F4 → C4F8
is 0.0469 M−1
· s
−1
. We start with 0.203 mol
C2F4 in a 2.00-liter container, with no C4F8
initially present. What will be the concentra-
tion of C2F4 after 1.00 hour ?
1
Expert's answer
2016-02-16T00:00:57-0500
For the second order reaction:
1/[A] = 1/[A]o + kt
But according to the provided equation, the process must be described as: 2A -----> products.
Initial concentration of C2F4 is: 0.203/2=0.1015 mol/l.
1/2·A = 1/ (2 · 0.1015) + 0.0469 · 3600 =173.766
2 A = 0.0058
A =0.0115 mol/l